OpenStudy (boxman61):

Solve: log(2)(7x-1)-log(2)(x-1)=4

OpenStudy (anonymous):

simple again log(a/b)= loga - logb

OpenStudy (saifoo.khan):

\[\log_2 ( 7x-1) - \log_2 (x-1) 4\log _2 2\]

OpenStudy (saifoo.khan):

MISSED THE = again! >_<

OpenStudy (boxman61):

\[\log_{2}(7x-1)-\log_{2}(x-1)=4 \]\[\log_{2}(7x-1/x+1)=4 \]\[7x-1/x+1=2^{4}\]\[7x-1=16(x+1)\]\[7x-1=16x+16\]\[-9x=17\]x=-17/9

OpenStudy (anonymous):

@Boxman61 yeahh ryt... u knw it very well..:)

OpenStudy (boxman61):

been awake for awhile so everythings blending together... thank you

OpenStudy (boxman61):

trying to prove... \[\log_{2}(7(-17/9)-1)-\log_{2}(-17/9+1)=4 \]\[\log_{2}(-128/9)-\log_{2}(-8/9)=4 \] im stuck there

OpenStudy (anonymous):

are apply abe formula only.. it will work... consider a= (-128/9) and b= (-8/9)

OpenStudy (anonymous):

apply the same formula*

OpenStudy (anonymous):

ohh here u mean there is no x to be find out.... so simply prove lhs=rhs...